# Tensor-Fused Multi-View Graph Contrastive Learning

### v1

@@ 1 Introduction
In recent years, graph neural networks (GNNs) have become indispensable tools for learning from graph-structured data prevalent in various domains such as social networks (Davies and Ajmeri,, 2022; Fan et al.,, 2019), molecular biology (Xia et al.,, 2023; Sun et al.,, 2021; Wang et al.,, 2022), and recommendation systems (Di Rocco et al.,, 2021; Yang et al.,, 2023). GNNs extend traditional neural networks by leveraging the relationships between nodes in a graph, using neighborhood aggregation to capture both node features and local graph structures. This method has proven effective in tasks such as node classification (Kipf and Welling,, 2016; Veličković et al.,, 2017), link prediction (Wang and Vinel,, 2021; Ai et al.,, 2022; Chen et al.,, 2022), and graph classification (Wen et al.,, 2024; Ying et al.,, 2018). However, GNNs often rely heavily on task-specific labels, which can be scarce or expensive to obtain in many real-world applications. This dependence on labeled data limits the generalization ability of GNNs, especially when applied to noisy or sparse graphs. To mitigate this issue, graph contrastive learning (GCL) has emerged as a promising solution in the field of unsupervised and semi-supervised learning (You et al., 2020a, ; Kefato and Girdzijauskas,, 2021). GCL enables models to learn meaningful representations by contrasting different augmented views of the same graph, thus reducing the need for extensive labeled data. By applying augmentation techniques such as node dropping, edge perturbation, and subgraph sampling, GCL encourages the model to capture invariant structural features, improving robustness and expressiveness in the face of data sparsity or noise.
@@
Despite the progress made by GCL, existing frameworks continue to face several challenges. A key limitation is their inability to capture higher-order structural relationships and latent topological information embedded in graphs. Most GCL models focus on local node-level or edge-level augmentations, which often overlook the global topological properties necessary for a complete understanding of graph data. This limitation is particularly evident in tasks where capturing complex, multi-scale topological features is crucial, such as molecular interactions or long-range dependencies in network systems. To overcome these issues, we propose the use of extended persistent homology (EPH), an advanced technique built on the foundations of persistent homology (PH) from topological data analysis (TDA). While PH has been studied by many articles (Horn et al.,, 2021; Wasserman,, 2018; Carlsson and Gabrielsson,, 2020), it is often limited to single filtrations, reducing its ability to capture multi-dimensional structures in complex graphs. EPH introduces multi-dimensional filtrations, offering a richer, more flexible representation of both local and global topological information, making it more effective for understanding complex topological structures. Several applications of EPH have shown promising results in tasks, including link prediction, node classification, and graph classification (Zhao et al.,, 2020; Yan et al.,, 2021; Carrière et al.,, 2020). Among them, TopoGCL (Chen et al., 2024e, ) is the first one to integrate EPH with GCL to capture important topological and geometric features and enhance representation learning. While TopoGCL shows notable promise in representation learning, it fails to (1) preserve the low-rank structure of representation tensors during learning, leading to inefficiencies and higher computational costs as the graph complexity increases, and (2) mitigate the additional noise introduced by EPH, which arises from its broader and more complex filtration process, potentially capturing irrelevant topological features and amplifying noise.
@@ 2 Related Work
Tensor-based Neural Networks. Tensors, which represent multidimensional data, are widely used across fields like neuroimaging (Zhou et al.,, 2013), economics (Chen et al.,, 2020; Liu and Chen,, 2022), international trade (Chen and Chen,, 2022), recommendation systems (Bi et al.,, 2018), and biomedical applications (Chen et al., 2024d, ; Chen et al., 2024a, ). Core areas of study include tensor decomposition (Zhang and Xia,, 2018; Chen et al., 2024a, ), tensor regression (Zhou et al.,, 2013; Li and Zhang,, 2017; Xia et al.,, 2022; Chen et al., 2024b, ), and tensor clustering (Sun and Li,, 2019; Mai et al.,, 2021; Luo and Zhang,, 2022). These methods have been effective in classification tasks, leading to techniques like support tensor machines (Hao et al.,, 2013; Guo et al.,, 2016), tensor discriminant analysis (Chen et al., 2024c, ), tensor logistic regression (Wimalawarne et al.,, 2016), and tensor neural networks (Kossaifi et al., 2020a, ; Wen et al.,, 2024).
While neural networks with tensor inputs efficiently process high-dimensional data, most approaches use tensors for computation rather than statistical analysis. Cohen et al., (2016) showed that deep networks can be understood through hierarchical tensor factorizations. Tensor contraction layers (Kossaifi et al.,, 2017) and regression layers (Kossaifi et al., 2020b, ) further regularize models, reducing parameters while maintaining accuracy. Recent advancements like the Graph Tensor Network (Xu et al.,, 2023) and Tensor-view Topological Graph Neural Network (Wen et al.,, 2024) offer new frameworks for deep learning on large, multi-dimensional data.
@@ 3 Preliminaries
Extended Persistent Homology. Persistent homology (PH) is a computational method in algebraic topology used to analyze the multi-scale topological features of data by tracking their evolution across different scales. Specifically, PH captures changes in topological properties, such as connected components, loops, and higher-dimensional holes, as a filtration parameter changes, allowing it to detect features that persist over a range of scales, which are indicative of significant structure in the data (Edelsbrunner et al.,, 2002; Zomorodian and Carlsson,, 2004). By using a multi-resolution approach, PH addresses the inherent limitations of classical homology and enables the extraction of latent shape properties of $\mathcal{G}$. The main approach involves selecting appropriate scale parameters $\gamma$ and analyzing how the shape of $\mathcal{G}$ evolves as $\mathcal{G}$ changes with respect to $\gamma$. Instead of studying $\mathcal{G}$ as a single object, we study a filtration $\mathcal{G}_{\gamma_{1}}\subseteq\ldots\subseteq\mathcal{G}_{\gamma_{n}}=\mathcal{G}$ where ${\gamma_{1}}<\cdots<{\gamma_{n}}$. We can then record the birth $b_{i}$ and death $d_{i}$ of each topological feature, representing the scale at which each feature first and last appears in the sublevel filtration. However, PH has limitations in retaining information about features that persist indefinitely, which results in the loss of crucial latent topological information for certain applications. To address these limitations, an alternative approach is to supplement the sublevel filtration with its superlevel counterpart, which is called extended persistent homology (EPH) (Chen et al., 2024e, ; Cohen-Steiner et al.,, 2009; Dey et al.,, 2014). Opposite to the sublevel filtration, we record the scale at which each feature first and last appears in the superlevel filtration $\mathcal{G}^{\gamma_{1}}\supseteq\mathcal{G}^{\gamma_{2}}\supseteq\ldots\supseteq\mathcal{G}^{\gamma_{n}}$. Hence, we can simultaneously assess topological features that persist and reappear in both directions, thereby providing a more comprehensive representation of the underlying topological properties of the data. Then, the extended topological information can be summarized as a multiset in $\mathbb{R}$, called the Extended Persistence Diagram (EPD), which can be denoted as $\text{EPD}=\{(b_{\rho},d_{\rho})\in\mathbb{R}^{2}\}$.
@@ 5.1 Experiment Settings
Baseline. We compare our proposed TensorMV-GCL on 11 real-world datasets with 15 state-of-the-art baselines including (1) Graphlet Kernel (GL) (Nino Shervashidze,, 2009), (2) Contrastive Multi-View Representation Learning on Graphs (MVGRL) (Hassani and Ahmadi,, 2020), (3) Weisfeiler-Lehman Sub-tree Kernel (WL) (Nino Shervashidze,, 2011), (4) Deep Graph Kernels (DGK) (Yanardag and Vishwanathan,, 2015), (5) node2vec  (Grover and Leskovec,, 2016), (6) sub2vec (Adhikari et al.,, 2017), (7) graph2vec  (Narayanan et al.,, 2017), (8) InfoGraph (Sun et al., 2019a, ), (9) Graph Contrastive Learning with Augmentations (GraphCL) (You et al., 2020b, ), (10) Graph Contrastive Learning Automated (JOAO) (You et al.,, 2021), (11) Adversarial Graph Augmentation to Improve Graph Contrastive Learning (AD-GCL) (Suresh et al.,, 2021), (12) AutoGCL (Yin et al.,, 2021), (13) A Review-aware Graph Contrastive Learning Framework for Recommendation (RGCL) (Shuai et al.,, 2022), (14) GCL-TAGS (Lin et al.,, 2022), and (15) TopoGCL (Chen et al., 2024e, ).

### v2

@@ 1 Introduction
In recent years, graph neural networks (GNNs) have become indispensable tools for learning from graph-structured data prevalent in various domains such as social networks (Davies and Ajmeri,, 2022; Fan et al.,, 2019), molecular biology (Xia et al.,, 2023; Sun et al.,, 2021; Wang et al.,, 2022), and recommendation systems (Di Rocco et al.,, 2021; Yang et al.,, 2023). GNNs extend traditional neural networks by leveraging the relationships between nodes in a graph, using neighborhood aggregation to capture both node features and local graph structures. This method has proven effective in tasks such as node classification (Kipf and Welling,, 2016; Veličković et al.,, 2017), link prediction (Wang and Vinel,, 2021; Ai et al.,, 2022; Chen et al., 2022b, ), and graph classification (Wen et al.,, 2024; Ying et al.,, 2018). However, GNNs often rely heavily on task-specific labels, which can be scarce or expensive to obtain in many real-world applications. This dependence on labeled data limits the generalization ability of GNNs, especially when applied to noisy or sparse graphs. To mitigate this issue, graph contrastive learning (GCL) has emerged as a promising solution in the field of unsupervised and semi-supervised learning (You et al., 2020a, ; Kefato and Girdzijauskas,, 2021). GCL enables models to learn meaningful representations by contrasting different augmented views of the same graph, thus reducing the need for extensive labeled data. By applying augmentation techniques such as node dropping, edge perturbation, and subgraph sampling, GCL encourages the model to capture invariant structural features, improving robustness and expressiveness in the face of data sparsity or noise.
@@
Despite the progress made by GCL, existing frameworks continue to face several challenges. A key limitation is their inability to capture higher-order structural relationships and latent topological information embedded in graphs. Most GCL models focus on local node-level or edge-level augmentations, which often overlook the global topological properties necessary for a complete understanding of graph data. This limitation is particularly evident in tasks where capturing complex, multi-scale topological features is crucial, such as molecular interactions or long-range dependencies in network systems. To overcome these issues, we propose the use of extended persistent homology (EPH), an advanced technique built on the foundations of persistent homology (PH) from topological data analysis (TDA). While PH has been studied by many articles (Horn et al.,, 2021; Wasserman,, 2018; Carlsson and Gabrielsson,, 2020), it is often limited to single filtrations, reducing its ability to capture multi-dimensional structures in complex graphs. EPH introduces multi-dimensional filtrations, offering a richer, more flexible representation of both local and global topological information, making it more effective for understanding complex topological structures. Several applications of EPH have shown promising results in tasks, including link prediction, node classification, and graph classification (Zhao et al.,, 2020; Yan et al.,, 2021; Carrière et al.,, 2020). Among them, TopoGCL (Chen et al., 2024j, ) is the first one to integrate EPH with GCL to capture important topological and geometric features and enhance representation learning. While TopoGCL shows notable promise in representation learning, it fails to (1) preserve the low-rank structure of representation tensors during learning, leading to inefficiencies and higher computational costs as the graph complexity increases, and (2) mitigate the additional noise introduced by EPH, which arises from its broader and more complex filtration process, potentially capturing irrelevant topological features and amplifying noise.
@@ 2 Related Work
Tensor-based Neural Networks. Tensor learning has seen rapid advancements across various domains, including statistical modeling, projection methods, and system identification Chen and Fan, (2023); Lin et al., (2021); Liu and Chen, (2022); Chen et al., 2024i ; Chen et al., 2024a ; Zhang et al., (2024); Li et al., (2023); Chen et al., 2024d ; Chen et al., 2024c . Notable advancements have emerged in areas such as community detection, constrained optimization, and transfer learning Chen et al., (2023, 2019, 2020); Chen and Chen, (2022); Chen et al., 2022a , as well as in reinforcement learning and distributed computing Chen et al., 2024h ; Chen et al., 2024e ; Chen et al., 2024b . Further developments in high-dimensional tensor classification, stochastic optimization, and statistical inference Chen et al., 2024f ; Chen et al., 2024g ; Chen et al., (2025); Xu et al., (2025) have expanded the applicability of tensor-based methodologies in modern machine learning.
While neural networks with tensor inputs efficiently process high-dimensional data, most approaches use tensors for computation rather than statistical analysis. Cohen et al., (2016) showed that deep networks can be understood through hierarchical tensor factorizations. Tensor contraction layers (Kossaifi et al.,, 2017) and regression layers (Kossaifi et al.,, 2020) further regularize models, reducing parameters while maintaining accuracy. Recent advancements like the Graph Tensor Network (Xu et al.,, 2023) and Tensor-view Topological Graph Neural Network (Wen et al.,, 2024) offer new frameworks for deep learning on large, multi-dimensional data. Additionally, substantial progress has been made in uncertainty quantification and representation learning using tensor-based approaches Wu et al., 2024a ; Wu et al., 2024b ; Wu et al., 2024c ; Albert et al., (2024).
@@ 3 Preliminaries
Extended Persistent Homology. Persistent homology (PH) is a computational method in algebraic topology used to analyze the multi-scale topological features of data by tracking their evolution across different scales. Specifically, PH captures changes in topological properties, such as connected components, loops, and higher-dimensional holes, as a filtration parameter changes, allowing it to detect features that persist over a range of scales, which are indicative of significant structure in the data (Edelsbrunner et al.,, 2002; Zomorodian and Carlsson,, 2004; Chen et al., 2022a, ). By using a multi-resolution approach, PH addresses the inherent limitations of classical homology and enables the extraction of latent shape properties of $\mathcal{G}$. The main approach involves selecting appropriate scale parameters $\gamma$ and analyzing how the shape of $\mathcal{G}$ evolves as $\mathcal{G}$ changes with respect to $\gamma$. Instead of studying $\mathcal{G}$ as a single object, we study a filtration $\mathcal{G}_{\gamma_{1}}\subseteq\ldots\subseteq\mathcal{G}_{\gamma_{n}}=\mathcal{G}$ where ${\gamma_{1}}<\cdots<{\gamma_{n}}$. We can then record the birth $b_{i}$ and death $d_{i}$ of each topological feature, representing the scale at which each feature first and last appears in the sublevel filtration. However, PH has limitations in retaining information about features that persist indefinitely, which results in the loss of crucial latent topological information for certain applications. To address these limitations, an alternative approach is to supplement the sublevel filtration with its superlevel counterpart, which is called extended persistent homology (EPH) (Chen et al., 2024j, ; Cohen-Steiner et al.,, 2009; Dey et al.,, 2014). Opposite to the sublevel filtration, we record the scale at which each feature first and last appears in the superlevel filtration $\mathcal{G}^{\gamma_{1}}\supseteq\mathcal{G}^{\gamma_{2}}\supseteq\ldots\supseteq\mathcal{G}^{\gamma_{n}}$. Hence, we can simultaneously assess topological features that persist and reappear in both directions, thereby providing a more comprehensive representation of the underlying topological properties of the data. Then, the extended topological information can be summarized as a multiset in $\mathbb{R}$, called the Extended Persistence Diagram (EPD), which can be denoted as $\text{EPD}=\{(b_{\rho},d_{\rho})\in\mathbb{R}^{2}\}$.
@@ 5.1 Experiment Settings
Baseline. We compare our proposed TensorMV-GCL on 11 real-world datasets with 15 state-of-the-art baselines including (1) Graphlet Kernel (GL) (Shervashidze et al.,, 2009), (2) Contrastive Multi-View Representation Learning on Graphs (MVGRL) (Hassani and Ahmadi,, 2020), (3) Weisfeiler-Lehman Sub-tree Kernel (WL) (Shervashidze et al.,, 2011), (4) Deep Graph Kernels (DGK) (Yanardag and Vishwanathan,, 2015), (5) node2vec  (Grover and Leskovec,, 2016), (6) sub2vec (Adhikari et al.,, 2017), (7) graph2vec  (Narayanan et al.,, 2017), (8) InfoGraph (Sun et al., 2019a, ), (9) Graph Contrastive Learning with Augmentations (GraphCL) (You et al., 2020b, ), (10) Graph Contrastive Learning Automated (JOAO) (You et al.,, 2021), (11) Adversarial Graph Augmentation to Improve Graph Contrastive Learning (AD-GCL) (Suresh et al.,, 2021), (12) AutoGCL (Yin et al.,, 2021), (13) A Review-aware Graph Contrastive Learning Framework for Recommendation (RGCL) (Shuai et al.,, 2022), (14) GCL-TAGS (Lin et al.,, 2022), and (15) TopoGCL (Chen et al., 2024j, ).